\begin{thebibliography}{10}

\bibitem{echo}
{Amazon Echo}.
\newblock
  https://www.amazon.com/Amazon-Echo-Bluetooth-Speaker-with-WiFi-Alexa/dp/B00X4WHP5E.

\bibitem{google}
{Google Home Features}.
\newblock \url{https://madeby.google.com/home/features/}.

\bibitem{bowden_icids16}
Kevin~K. Bowden, Grace~I. Lin, Lena~I. Reed, Jean E.~Fox Tree, and Marilyn~A.
  Walker.
\newblock {M2D:} monolog to dialog generation for conversational story telling.
\newblock In {\em Interactive Storytelling - 9th International Conference on
  Interactive Digital Storytelling, {ICIDS} 2016, Los Angeles, CA, USA,
  November 15-18, 2016, Proceedings}, pages 12--24, 2016.

\bibitem{Henderson2014}
Matthew Henderson, Blaise Thomson, and Jason Williams.
\newblock The second dialog state tracking challenge.
\newblock In {\em Proceedings of SIGDIAL}. ACL – Association for
  Computational Linguistics, June 2014.

\bibitem{Higashinaka14}
Ryuichiro Higashinaka, Kenji Imamura, Toyomi Meguro, Chiaki Miyazaki, Nozomi
  Kobayashi, Hiroaki Sugiyama, Toru Hirano, Toshiro Makino, and Yoshihiro
  Matsuo.
\newblock Towards an open-domain conversational system fully based on natural
  language processing.
\newblock In {\em COLING}. ACL, 2014.

\bibitem{hirschman2000evaluating}
Lynette Hirschman.
\newblock Evaluating spoken language interaction: Experiences from the darpa
  spoken language program 1990--1995.
\newblock {\em Spoken Language Discourse. MIT Press, Cambridge, Mass}, 2000.

\bibitem{DBLP:journals/corr/KadlecSK15}
Rudolf Kadlec, Martin Schmid, and Jan Kleindienst.
\newblock Improved deep learning baselines for ubuntu corpus dialogs.
\newblock {\em CoRR}, abs/1510.03753, 2015.

\bibitem{li2016persona}
Jiwei Li, Michel Galley, Chris Brockett, Jianfeng Gao, and Bill Dolan.
\newblock {A Persona-Based Neural Conversation Model}.
\newblock {\em arXiv preprint arXiv:1603.06155}, 2016.

\bibitem{DBLP:conf/sigdial/LowePSP15}
Ryan Lowe, Nissan Pow, Iulian Serban, and Joelle Pineau.
\newblock The ubuntu dialogue corpus: {A} large dataset for research in
  unstructured multi-turn dialogue systems.
\newblock In {\em Proceedings of the {SIGDIAL} 2015 Conference, The 16th Annual
  Meeting of the Special Interest Group on Discourse and Dialogue, 2-4
  September 2015, Prague, Czech Republic}, pages 285--294, 2015.

\bibitem{DBLP:journals/dad/LowePSCLP17}
Ryan~Thomas Lowe, Nissan Pow, Iulian~Vlad Serban, Laurent Charlin, Chia{-}Wei
  Liu, and Joelle Pineau.
\newblock Training end-to-end dialogue systems with the ubuntu dialogue corpus.
\newblock {\em D{\&}D}, 8(1):31--65, 2017.

\bibitem{lukin_games14}
Stephanie Lukin, James~Owen Ryan, and Marilyn~A Walker.
\newblock Automating direct speech variations in stories and games.
\newblock In {\em Proceedings of the 3rd Workshop on Games and NLP}, October
  2014.

\bibitem{Misraetal_naacl15}
Amita Misra, Pranav Anand, Jean E.~Fox Tree, and Marilyn~A. Walker.
\newblock Using summarization to discover argument facets in online idealogical
  dialog.
\newblock In {\em {NAACL} {HLT} 2015, The 2015 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Denver, Colorado, USA, May 31 - June 5, 2015}, 2015.

\bibitem{Misraetal_sigdial16}
Amita Misra, Brian Ecker, and Marilyn~A. Walker.
\newblock Measuring the similarity of sentential arguments in dialogue.
\newblock In {\em Proceedings of the {SIGDIAL} 2016 Conference, The 17th Annual
  Meeting of the Special Interest Group on Discourse and Dialogue, 13-15
  September 2016, Los Angeles, CA, {USA}}, 2016.

\bibitem{oraby_sigdial16}
Shereen Oraby, Vrindavan Harrison, Lena Reed, Ernesto Hernandez, Ellen Riloff,
  and Marilyn~A. Walker.
\newblock Creating and characterizing a diverse corpus of sarcasm in dialogue.
\newblock In {\em Proceedings of the {SIGDIAL} 2016 Conference, The 17th Annual
  Meeting of the Special Interest Group on Discourse and Dialogue, 13-15
  September 2016, Los Angeles, CA, {USA}}, pages 31--41, 2016.

\bibitem{oraby_am15}
Shereen Oraby, Lena Reed, Ryan Compton, Ellen Riloff, Marilyn~A. Walker, and
  Steve Whittaker.
\newblock And that's {A} fact: Distinguishing factual and emotional
  argumentation in online dialogue.
\newblock In {\em Proceedings of the 2nd Workshop on Argumentation Mining,
  ArgMining@HLT-NAACL 2015, June 4, 2015, Denver, Colorado, {USA}}, pages
  116--126, 2015.

\bibitem{price1992subject}
Patti Price, Lynette Hirschman, Elizabeth Shriberg, and Elizabeth Wade.
\newblock Subject-based evaluation measures for interactive spoken language
  systems.
\newblock In {\em Proceedings of the workshop on Speech and Natural Language},
  pages 34--39. Association for Computational Linguistics, 1992.

\bibitem{Ritter2011}
Alan Ritter, Colin Cherry, and William~B. Dolan.
\newblock Data-driven response generation in social media.
\newblock In {\em Proceedings of the Conference on Empirical Methods in Natural
  Language Processing}. Association for Computational Linguistics, 2011.

\bibitem{Serban2015}
Iulian~Vlad Serban, Ryan Lowe, Laurent Charlin, and Joelle Pineau.
\newblock A survey of available corpora for building data-driven dialogue
  systems.
\newblock {\em CoRR}, abs/1512.05742, 2015.

\bibitem{sordoni2015neural}
Alessandro Sordoni, Michel Galley, Michael Auli, Chris Brockett, Yangfeng Ji,
  Margaret Mitchell, Jian-Yun Nie, Jianfeng Gao, and Bill Dolan.
\newblock {A Neural Network Approach to Context-Sensitive Generation of
  Conversational Responses}.
\newblock {\em arXiv preprint arXiv:1506.06714}, 2015.

\bibitem{swanson2008say}
Reid Swanson and Andrew~S Gordon.
\newblock {Say Anything: A Massively Collaborative Open Domain Story Writing
  Companion}.
\newblock In {\em Interactive Storytelling}, pages 32--40. Springer, 2008.

\bibitem{Tsiakoulis12}
Pirros Tsiakoulis, Milica Gasic, Matthew Henderson, Joaquin Plannels, Jorge
  Prombonas, Blaise Thomson, Kai Yu, Steve Young, and Eli Tzirkel.
\newblock Statistical methods for building robust spoken dialogue systems in an
  automobile.
\newblock In {\em in 4th International Conference on Applied Human Factors and
  Ergonomics}, 2012.

\bibitem{vinyals2015neural}
Oriol Vinyals and Quoc Le.
\newblock {A Neural Conversational Model}.
\newblock {\em arXiv preprint arXiv:1506.05869}, 2015.

\bibitem{walker1997paradise}
Marilyn~A Walker, Diane~J Litman, Candace~A Kamm, and Alicia Abella.
\newblock Paradise: A framework for evaluating spoken dialogue agents.
\newblock In {\em Proceedings of the eighth conference on European chapter of
  the Association for Computational Linguistics}, pages 271--280. Association
  for Computational Linguistics, 1997.

\bibitem{walker2001quantitative}
Marilyn~A Walker, Rebecca Passonneau, and Julie~E Boland.
\newblock Quantitative and qualitative evaluation of darpa communicator spoken
  dialogue systems.
\newblock In {\em Proceedings of the 39th Annual Meeting on Association for
  Computational Linguistics}, pages 515--522. Association for Computational
  Linguistics, 2001.

\end{thebibliography}

\documentclass[graybox]{svmult}

\usepackage{mathptmx}       
\usepackage{helvet}         
\usepackage{courier}        
\usepackage{type1cm}        
\usepackage{makeidx}         
\usepackage{graphicx}        
\usepackage{multicol}        
\usepackage[bottom]{footmisc}

\makeindex             

\begin{document}

\title*{Data-Driven Dialogue Systems for Social Agents}
\titlerunning{Data-Driven Dialogue Systems for Social Agents}
\author{{\bf Kevin K. Bowden, Shereen Oraby, Amita Misra, Jiaqi Wu, and Stephanie Lukin} \\
Natural Language and Dialogue Systems Lab \\ 
University of California, Santa Cruz \\
\{kkbowden, soraby, amisra2, jwu64, slukin\}@ucsc.edu}
\authorrunning{K. Bowden, S. Oraby, A. Misra, J. Wu, and S. Lukin} 

\date{}

%
%

\maketitle

\abstract{In
  order to build dialogue systems to tackle the ambitious task of
  holding social conversations, we argue that we need a data-driven approach that includes insight into human conversational ``chit-chat'', and which
  incorporates different natural language processing modules. Our
  strategy is to analyze and index large corpora of social media data,
  including Twitter conversations, online debates, dialogues
  between friends, and blog posts, and then to couple this data
  retrieval with modules that perform tasks such as sentiment and
  style analysis, topic modeling, and summarization. We aim for
  personal assistants that can learn more nuanced human language, and
  to grow from task-oriented agents to more personable social
  bots.  }
  
\section{From Task-Oriented Agents to Social Bots}
Devices like the Amazon Echo and Google Home have entered our homes to perform task-oriented functions,
such as looking up today's headlines and setting reminders
\cite{echo,google}. As these devices evolve, we have begun to expect social conversation, where the device must learn to personalize and produce natural language style.

Social conversation is not explicitly goal-driven in the same way as task-oriented dialogue. Many dialogue systems in both the written and spoken medium have been developed for task-oriented agents with an explicit goal of restaurant information retrieval, booking a flight, diagnosing an IT issue, or providing automotive customer support \cite{hirschman2000evaluating,price1992subject,walker1997paradise,walker2001quantitative,Henderson2014,Tsiakoulis12}. 
These tasks often revolve around question answering, with little ``chit-chat''.
Templates are often used for generation and state tracking, but since
they are optimized for the task at hand, the conversation can 
either become stale, or maintaining a conversation requires the 
intractable task of  manually authoring many different social interactions
 that can be used in a particular context.

We argue that a social agent should be
spontaneous, and allow for human-friendly conversations that do not follow a perfectly-defined trajectory. In
order to build such a conversational dialogue system, we exploit the abundance of human-human social media
conversations, and develop methods informed by natural language
processing modules that model, analyze, and generate utterances that better suit the context. 

\section{Data-Driven Models of Human Language}

A myriad of social media data has led to the development of new techniques
for language understanding from open domain conversations, and many
corpora are available for building data-driven dialogue systems
\cite{Ritter2011,Serban2015}.
%
While there are differences between how people speak in person and in an online text-based environment, the social agents we build should not be limited in their language; they should be exposed to many different styles and vocabularies. Online conversations can be repurposed in new dialogues, but only if they can be properly indexed or adapted to the context. Data retrieval algorithms have been successfully employed to
co-construct an unfolding narrative between the user and computer
\cite{swanson2008say}, and re-use existing conversations \cite{Higashinaka14}. Other approaches train on such
conversations to analyze sequence and word patterns, but lack detailed
annotations and analysis, such as emotion and humor
\cite{sordoni2015neural,vinyals2015neural,li2016persona}. The large Ubuntu Dialogue Corpus \cite{DBLP:conf/sigdial/LowePSP15} with over 7 million utterances is large enough to train neural network models \cite{DBLP:journals/corr/KadlecSK15, DBLP:journals/dad/LowePSCLP17}.

We argue that combining data-driven retrieval with modules for
sentiment analysis and style, topic analysis, summarization,
paraphrasing, and rephrasing will allow for more human-like social
conversation. This requires that data be indexed based on domain and
requirement, and then retrieve candidate utterances based on dialogue
state and context. Likewise, in order to avoid stale and repetitive utterances, we
can alter and repurpose the candidate utterances; for
example, we can use paraphrase or summarization to create new ways of
saying the same thing, or to select utterance candidates
according to the desired sentiment
\cite{Misraetal_naacl15,Misraetal_sigdial16}. The style of
an utterance can be altered based on requirements; introducing
elements of sarcasm, or aspects of factual and emotional argumentation
styles \cite{oraby_am15,oraby_sigdial16}. Changes in the perceived
speaker personality can also make more
personable conversations \cite{lukin_games14}. Even utterances from
monologic texts can be leveraged by converting the content to dialogic
flow, and performing stylistic
transformations \cite{bowden_icids16}.

Of course, while many data sources may be of interest for indexing knowledge
for a dialogue system, annotations are not always available or easy to obtain.
By using machine learning models designed to classify different classes of interest,
such as sentiment, sarcasm, and topic, data can be bootstrapped to greatly increase
the amount of data available for indexing and utterance selection \cite{oraby_am15}.

There is no shortage of human generated dialogues, but the challenge is
to analyze and harness them appropriately for social-dialogue
generation. We aim to combine data-driven methods to repurpose
existing social media dialogues, in addition to a suite of tools for
sentiment analysis, topic identification, summarization, paraphrase,
and rephrasing, to  develop a socially-adept
agent that can carry on a natural conversation.

\bibliographystyle{plain}
\bibliography{author}

\end{document}